# On the transitivity of Anosov diffeomorphisms

### v2

@@ 7.3.1. The $C^{1}$ interior of $\mathcal{WIP}_{h}$
()($\subseteq$)( ⊆ ) Take $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$. By Lemma 7.11, $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})$, and, by Proposition 7.4, $f$ is $\Omega-$stable. Since $f$ also has the weak $su-$intersecting property in $\operatorname{Per}(f)$, then $f$ is chain-transitive by Theorem 6.1. So, $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$.
()superset-of-or-equals($\supseteq$)( ⊇ ) Suppose that $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$. By Corollary 7.6, $f$ is a topologically transitive Anosov diffeomorphism. As the space of topologically transitive Anosov diffeomorphisms is open, there exists a $C^{1}$ open neighborhood $\mathcal{U}$ of $f$ such that any $g\in\mathcal{U}$ is a topologically transitive Anosov diffeomorphism. By Theorem 6.1, every $g\in\mathcal{U}$ has the weak $su-$intersecting property in $\operatorname{Per}(g)$. So, $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$.
We are ready to complete the proof of Proposition 7.10 since, by Corollary 7.6, a diffeomorphism $f$ belongs in $\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})=\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$ if and only if it is a topologically transitive Anosov diffeomorphism. ∎
Consequently:
@@
Firstly, by (6.1), [53][Lemma 2.1] and Lemma 6.4, one has
@@
Now, take $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$ and $\mathcal{U}$ a $C^{1}$ open neighborhood of $f$ contained in $\mathcal{WIP}_{h}$. By Proposition 7.10, every $g\in\mathcal{U}$ is a topologically transitive Anosov diffeomorphism. In particular, every $g\in\mathcal{U}$ is $\Omega-$stable and has the weak $su-$intersecting property in $\operatorname{Per}(g)$. Thus, by Theorem 6.1, every $g\in\mathcal{U}$ has the two-sided limit shadowing property, the specification property and the gluing orbit property. Hence,
@@ 7.3.2. Generic elements of $\mathcal{WIP}_{h}$
In [14][Theorem C], B. Carvalho showed that $C^{1}$ generically if a diffeomorphism $f\colon M\to M$ satisfies the two-sided limit shadowing property then $f$ is a topologically transitive Anosov diffeomorphism. In [36][Corollary 1.2], K. Lee and K. Tajbakhsh established that $C^{1}$ generically if $f$ has the specification property then $f$ is a topologically transitive Anosov diffeomorphism. In [34], M. Lee proved that $C^{1}$ generically if $f$ has the barycenter property in $\operatorname{Per}(f)$ then it is $\Omega-$stable. We improve these results as follows. Recall from (7.1) that $\mathcal{R}=\mathcal{KS}\cap\mathcal{I}$.
@@
If $f\in\mathcal{R}\,\cap\,\mathcal{WIP}_{h}$, then $f$ is a topologically transitive Anosov diffeomorphism.
@@
()($\subseteq$)( ⊆ ) Take $f\in\mathcal{R}\,\cap\,\mathcal{WIP}_{h}$, that is, $f\in\mathcal{KS}$ and $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. By Corollary 7.3, all periodic points of $f$ are hyperbolic and have the same index. Thus, $f\in\mathcal{PCI}$. By Proposition 7.7, $f$ is $\Omega-$stable. Since, in addition, $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, from Theorem 6.1 we deduce that $f$ is chain-transitive. Hence, $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\,\cap\,\mathcal{CT}$.
()superset-of-or-equals($\supseteq$)( ⊇ ) Suppose that $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\cap\mathcal{CT}$. By Proposition 7.7, $f$ is $\Omega-$ stable. Thus, $f$ is chain-transitive. Therefore, by Theorem 6.1, $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. ∎
Let us complete the proof of Proposition 7.13. Suppose that $f\in\mathcal{R}\,\cap\,\mathcal{WIP}_{h}$. By Lemma 7.14, $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\,\cap\,\mathcal{CT}$. Consequently, by Proposition 7.7, $f$ is $\Omega-$ stable. Since $f$ is also chain-transitive, Theorem 6.1 ensures that $f$ is a topologically transitive Anosov diffeomorphism. ∎
The next corollary should be compared with Theorem 6.1.
@@
$(6)\Rightarrow(1)$ Let $f$ be a topologically transitive Anosov diffeomorphism. So $f$ is an expansive topologically transitive diffeomorphism that has the shadowing property. Since $M$ is connected, $f$ is topologically mixing. By [30, Theorem 45], $f$ has the specification property. Thus, according to
Theorem 5.1, $f$ has the two-sided limit shadowing property.
$(6)\Rightarrow(7)$ This was established in the implication $(6)\Rightarrow(7)$ of Theorem 6.1.
$(7)\Rightarrow(4)$ Assume that $f$ has the specification property. By [53][Lemma 2.1], we know that $f$ has the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$.
@@ 7.4. Application 4
The proof of Theorem 6.1 ensures, without the assumption of $\Omega-$stability, that:
the specification property $\Rightarrow$ the almost specification property
$\Rightarrow$ the asymptotic average shadowing property
$\Rightarrow$ the average shadowing property
@@
As regards these properties, in [52], K. Sakai showed that the $C^{1}$ interior of the set of diffeomorphisms of a compact surface satisfying the average shadowing property is equal to the set of topologically transitive Anosov diffeomorphisms. On the other hand, in [40], R. Mañé proved that every $C^{1}-$robustly transitive diffeomorphism of a compact surface is a topologically transitive Anosov diffeomorphism. As a consequence of Theorem 6.1 together with Corollary 7.2, we provide a unique proof for both these results.
@@
Let $H$ be a chain-transitive diffeomorphism. Then $H^{-1}$ is also chain-transitive (cf. [64]). Moreover, $H$ has no sinks (cf. [32][Lemma 2.1]). Since a source of $H$ is a sink of $H^{-1}$, we deduce that $H$ has no sources as well. ∎
Let us resume the proof of Proposition 7.16. As $\tilde{h}$ is chain-transitive, by Lemma 7.18 we know that $\tilde{h}$ does not have sinks or sources. Therefore, as $\tilde{h}\in\mathcal{Z}$, Lemma 7.17 yields that $\tilde{h}$ is $\Omega-$stable. Thus Theorem 6.1 ensures that $\tilde{h}$ is a topologically transitive Anosov diffeomorphism. Hence, by Corollary 7.2, for any pair of hyperbolic periodic points $p$ and $q$ of $\tilde{h}$ one has $\operatorname{ind}(p)=\operatorname{ind}(q)$. Anyway, this contradicts the fact that $\tilde{h}$ has two hyperbolic periodic points $\tilde{p}_{1},\tilde{p}_{2}$ with different indices. Consequently, $f$ is $\Omega-$stable. Since $f$ is also chain-transitive, Theorem 6.1 indicates that $f$ is a topologically transitive Anosov diffeomorphism. This ends the proof of Proposition 7.16. ∎
@@
When the dimension of the manifold is equal to $3$ or $4$ it is well-known that a $C^{1}-$robustly transitive diffeomorphism may not be an Anosov diffeomorphism [38, 54]. In [8], A. Arbieto and R. Ribeiro showed that the $C^{1}$ interior of the set of flows with the asymptotic average shadowing property (or with the average shadowing property) on a $3-$dimensional smooth closed manifold is formed by topologically transitive Anosov flows.
@@ 7.5. Application 5
$(b)\Rightarrow(c)$ This was established in [7][Proposition 3.1].
$(c)\Rightarrow(a)$ Take $f\in{\rm Diff}^{1}_{m}(M)$ complying with the condition $(c)$. We claim that $f$ is an Anosov diffeomorphism. Assume otherwise. By [7][Theorem 1.1], $f\notin\mathcal{F}^{1}_{m}(M)$, so there exists $g\in{\rm Diff}^{1}_{m}(M)$ close enough to $f$ with a non-hyperbolic periodic point $p$. Thus, according to [7][Lemma 5.1], there exists $h\in{\rm Diff}^{1}_{m}(M)$ which is close enough to $f$ and has two periodic points $q$ and $r$ with different indices. This contradicts $(c)$. ∎
@@
Denote by $\mathcal{KS}_{m}$ the $C^{1}$ generic set of Kupka-Smale volume-preserving diffeomorphisms provided by Robinson’s theorem (see [51]). The coming lemma is a version in the conservative setting of Corollary 7.3, and is proved by a similar argument.
@@
$(3)\Rightarrow(4)$ Suppose that, for any $\,p,q\in\operatorname{Per}_{h}(f)$, we have $\operatorname{ind}(p)=\operatorname{ind}(q)$. We claim that $f\in\mathcal{F}^{1}_{m}(M)$. Assume otherwise, that $f\notin\mathcal{F}^{1}_{m}(M)$, so there exists $g\in{\rm Diff}^{1}_{m}(M)$ close enough to $f$ with a non-hyperbolic periodic point $p$. Thus, according to [7][Lemma 5.1], there exists $h\in{\rm Diff}^{1}_{m}(M)$ which is close enough to $f$ and has two periodic points with different indices. Since $f\in\mathcal{J}$, by Lemma 7.22, $f$ has two hyperbolic periodic points with different indices. Yet, this contradicts the assumption on $f$. Therefore, $f\in\mathcal{F}^{1}_{m}(M)$.
$(4)\Rightarrow(5)$ This was established in [7][Theorem 1.1].
$(5)\Rightarrow(1)$ This is clear by Theorem A(a). ∎
7.6. Application 6
Given a smooth closed Riemannian manifold $M$, denote by $\mathbb{O}(M)$ the space of $\Omega-$stable $C^{1}$ diffeomorphisms of $M$ and by $\mathbb{TA}(M)$ its subspace of topologically transitive Anosov diffeomorphisms. The following consequences of Proposition 7.7 convey an insight into the closure of $\mathbb{O}(M)$.
Proposition 7.24.
Let $M$ be a smooth closed Riemannian manifold with dimension $\geqslant 2$. There is a $C^{1}$ Baire generic set $\mathcal{S}\subset\operatorname{Diff}^{1}(M)$ such that, if $f\in\mathcal{S}\cap\operatorname{int}_{C^{1}}\Big{(}\overline{\mathbb{O}(M)}\setminus\overline{\mathbb{TA}(M)}\Big{)}$, then $f\in\mathbb{O}(M)\setminus\overline{\mathbb{TA}(M)}$.
Proof.
Observe that $\mathbb{O}(M)\setminus\mathbb{TA}(M)\neq\emptyset$, since Morse-Smale diffeomorphisms exist on all smooth closed Riemannian manifolds (cf. [56]).
Claim: $\mathbb{O}(M)\setminus\mathbb{TA}(M)=\mathbb{O}(M)\setminus\overline{\mathbb{TA}(M)}$.
Proof.
Clearly, $\mathbb{O}(M)\setminus\overline{\mathbb{TA}(M)}\subseteq\mathbb{O}(M)\setminus\mathbb{TA}(M)$. On the other hand, assume that there exists $f\in\mathbb{O}(M)\setminus\mathbb{TA}(M)$ such that $f\in\overline{\mathbb{TA}(M)}$. Since the $\Omega-$stability is an open condition and $f\in\overline{\mathbb{TA}(M)}$, then $f_{|\Omega(f)}$ is conjugated to the restriction of a topologically transitive Anosov diffeomorphism to its non-wandering set (which is $M$). Therefore, $f$ is topologically transitive and, since $f$ is $\Omega-$stable, $\Omega(f)=M$ is hyperbolic. Thus, $f\in\mathbb{TA}$, contradicting the initial assumption on $f$. ∎
Note that, from the previous Claim we conclude that $\mathbb{O}(M)\setminus\mathbb{TA}(M)$ is an open set of $\operatorname{Diff}^{1}(M)$. Thus, $\operatorname{int}_{C^{1}}\big{(}\overline{\mathbb{O}(M)}\setminus\overline{\mathbb{TA}(M)}\big{)}\neq\emptyset$.
Lemma 7.25.
Let $Y\subset\operatorname{Diff}^{1}(M)$ an open set. There exists a $C^{1}$ open and dense subset $\mathcal{S}$ of ${\rm Diff}^{1}(M)$ such that, if $f\in\mathcal{S}\cap\overline{Y}$, then $f\in Y$.
Proof.
Consider the complement $Z={\rm Diff}^{1}(M)\setminus\overline{Y}$. Hence, the set $\mathcal{S}=Y\cup Z$ is open in ${\rm Diff}^{1}(M)$. We claim that $\mathcal{S}$ is also dense in ${\rm Diff}^{1}(M)$. Indeed, if $g\in{\rm Diff}^{1}(M)\setminus Z$, then $g\in\overline{Y}$, and so it is approximated by diffeomorphisms of $Y\subseteq\mathcal{S}$. Moreover,
$f\in\mathcal{S}\cap\overline{Y}\quad\Rightarrow\quad f\in(Y\cup Z)\cap\overline{Y}\quad\Rightarrow\quad f\in Y.$
Applying Lemma 7.25 to $Y=\mathbb{O}(M)$, we find a $C^{1}$ Baire generic set $\mathcal{S}\subset{\rm Diff}^{1}(M)$ such that if $f\in\mathcal{S}\,\cap\,\overline{\mathbb{O}(M)}$ then $f\in\mathbb{O}(M)$. Therefore, $\mathcal{S}\cap\operatorname{int}_{C^{1}}\big{(}\overline{\mathbb{O}(M)}\setminus\overline{\mathbb{TA}(M)}\big{)}\neq\emptyset$. Moreover, if $f\in\mathcal{S}\cap\big{(}\operatorname{int}_{C^{1}}\big{(}\overline{\mathbb{O}(M)}\setminus\overline{\mathbb{TA}(M)}\big{)}\big{)}$, then $f\in\big{(}\mathcal{S}\cap\overline{\mathbb{O}(M)}\big{)}\setminus\overline{\mathbb{TA}(M)}$. Hence, by Lemma 7.25, $f\in\mathbb{O}(M)\setminus\overline{\mathbb{TA}(M)}$. The proof of Proposition 7.24 is complete. ∎
Recall from (7.1) that $\mathcal{R}=\mathcal{KS}\cap\mathcal{I}$.
Corollary 7.26.
If $f\in\mathcal{R}\cap\overline{\mathbb{TA}(M)}$, then $f$ is a topologically transitive Anosov diffeomorphism.
Proof.
We start by showing that the property of having a constant index within the hyperbolic periodic points extends to the $C^{1}$ closure of the space $\mathbb{TA}(M)$.
Lemma 7.27.
If $g\in\overline{\mathbb{TA}(M)}$, then $g\in\mathcal{PCI}$.
Proof.
Clearly, if $\operatorname{Per}_{h}(g)=\emptyset$, then $g\in\mathcal{PCI}.$ Assume now that there is $g\in\overline{\mathbb{TA}(M)}$ such that $\operatorname{Per}_{h}(g)\neq\emptyset$ but $g\notin\mathcal{PCI}$. Let $\mathcal{N}={\rm Diff}^{1}(M)\setminus\mathcal{PCI}$ be the set of $C^{1}$ diffeomorphisms of $M$ which have two hyperbolic periodic points with different indices. The hyperbolicity of these periodic points guarantees that $\mathcal{N}$ is open in ${\rm Diff}^{1}(M)$. So, $g\in\mathcal{N}$ and, given that $g$ is in $\overline{\mathbb{TA}(M)}$, there exists $h\in\mathbb{TA}(M)$ such that $h\in\mathcal{N}$. However, the existence of such an $h$ in $\mathcal{N}$ contradicts Corollary 7.2. Consequently, $g\in\mathcal{PCI}$. ∎
To complete the proof of Corollary 7.26, take $f\in\mathcal{R}\cap\overline{\mathbb{TA}(M)}=\mathcal{KS}\cap\mathcal{I}\cap\overline{\mathbb{TA}(M)}$. By Lemma 7.27, we know that $f\in\mathcal{PCI}$. So, $f\in\mathcal{R}\cap\mathcal{PCI}$, and therefore $f$ is $\Omega-$stable by Proposition 7.7. Since the $\Omega$-stability is an open condition and $f\in\overline{\mathbb{TA}(M)}$, there exists $h\in\mathbb{TA}(M)$ such that $f$ and $h$ are conjugate. Thus, $f$ is $\Omega-$stable and topologically transitive; hence, $f$ is an Anosov diffeomorphism. ∎

### v3

@@ 7.3.1. The $C^{1}$ interior of $\mathcal{WIP}_{h}$
()($\subseteq$)( ⊆ ) Take $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$. By Lemma 7.11, $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})$, and, by Proposition 7.4, $f$ is $\Omega-$stable. Since $f$ also has the weak $su-$intersecting property in $\operatorname{Per}(f)$, then $f$ is chain-transitive by Theorem C. So, $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$.
()superset-of-or-equals($\supseteq$)( ⊇ ) Suppose that $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$. By Corollary 7.6, $f$ is a topologically transitive Anosov diffeomorphism. As the space of topologically transitive Anosov diffeomorphisms is open, there exists a $C^{1}$ open neighborhood $\mathcal{U}$ of $f$ such that any $g\in\mathcal{U}$ is a topologically transitive Anosov diffeomorphism. By Theorem C, every $g\in\mathcal{U}$ has the weak $su-$intersecting property in $\operatorname{Per}(g)$. So, $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$.
We are ready to finalize the proof of Proposition 7.10 since, by Corollary 7.6, a diffeomorphism $f$ belongs in $\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})=\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$ if and only if it is a topologically transitive Anosov diffeomorphism. ∎
@@
Firstly, by (6.1), [53, Lemma 2.1] and Lemma 6.2, one has
@@
Now, take $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$ and $\mathcal{U}$ a $C^{1}$ open neighborhood of $f$ contained in $\mathcal{WIP}_{h}$. By Proposition 7.10, every $g\in\mathcal{U}$ is a topologically transitive Anosov diffeomorphism. In particular, every $g\in\mathcal{U}$ is $\Omega-$stable and has the weak $su-$intersecting property in $\operatorname{Per}(g)$. Thus, by Theorem C, every $g\in\mathcal{U}$ has the two-sided limit shadowing property, the specification property and the gluing orbit property. Hence,
@@ 7.3.2. Generic elements of $\mathcal{WIP}_{h}$
In [14, Theorem C], B. Carvalho showed that $C^{1}$ generically if a diffeomorphism $f\colon M\to M$ satisfies the two-sided limit shadowing property then $f$ is a topologically transitive Anosov diffeomorphism. In [36, Corollary 1.2], K. Lee and K. Tajbakhsh established that $C^{1}$ generically if $f$ has the specification property then $f$ is a topologically transitive Anosov diffeomorphism. In [34], M. Lee proved that $C^{1}$ generically if $f$ has the barycenter property in $\operatorname{Per}(f)$ then it is $\Omega-$stable. We improve these results as follows.
@@
Let $\mathcal{R}=\mathcal{KS}\cap\mathcal{I}$ be the set defined in (7.1). If $f\in\mathcal{R}\,\cap\,\mathcal{WIP}_{h}$, then $f$ is a topologically transitive Anosov diffeomorphism.
@@
()($\subseteq$)( ⊆ ) Take $f\in\mathcal{R}\,\cap\,\mathcal{WIP}_{h}$, that is, $f\in\mathcal{KS}$ and $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. By Corollary 7.3, all periodic points of $f$ are hyperbolic and have the same index. Thus, $f\in\mathcal{PCI}$. By Proposition 7.7, $f$ is $\Omega-$stable. Since, in addition, $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, from Theorem C we deduce that $f$ is chain-transitive. Hence, $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\,\cap\,\mathcal{CT}$.
()superset-of-or-equals($\supseteq$)( ⊇ ) Suppose that $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\cap\mathcal{CT}$. By Proposition 7.7, $f$ is $\Omega-$ stable. Thus, $f$ is chain-transitive. Therefore, by Theorem C, $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. ∎
Let us end the proof of Proposition 7.13. Suppose that $f\in\mathcal{R}\,\cap\,\mathcal{WIP}_{h}$. By Lemma 7.14, $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\,\cap\,\mathcal{CT}$. Consequently, by Proposition 7.7, $f$ is $\Omega-$ stable. Since $f$ is also chain-transitive, Theorem C ensures that $f$ is a topologically transitive Anosov diffeomorphism. ∎
@@
$(6)\Rightarrow(1)$ Let $f$ be a topologically transitive Anosov diffeomorphism. So $f$ is an expansive topologically transitive diffeomorphism that has the shadowing property. Since $M$ is connected, $f$ is topologically mixing. By [30, Theorem 45], $f$ has the specification property. Thus, according to Proposition 5.1, $f$ has the two-sided limit shadowing property.
$(6)\Rightarrow(7)$ This was established in the implication $(6)\Rightarrow(7)$ of Theorem C.
$(7)\Rightarrow(4)$ Assume that $f$ has the specification property. By [53, Lemma 2.1], we know that $f$ has the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$.
@@ 7.4. Application 4
The proof of Theorem C ensures, without the assumption of $\Omega-$stability, that
specification property $\Rightarrow$ almost specification property
$\Rightarrow$ asymptotic average shadowing property
$\Rightarrow$ average shadowing property
@@
As regards these properties, in [52], K. Sakai showed that the $C^{1}$ interior of the set of diffeomorphisms of a compact surface satisfying the average shadowing property is equal to the set of topologically transitive Anosov diffeomorphisms. On the other hand, in [40], R. Mañé proved that every $C^{1}-$robustly transitive diffeomorphism of a compact surface is a topologically transitive Anosov diffeomorphism. As a consequence of Theorem C together with Corollary 7.2, we provide a unique proof for both these results.
@@
Let $H$ be a chain-transitive diffeomorphism. Then $H^{-1}$ is also chain-transitive (cf. [63]). Moreover, $H$ has no sinks (cf. [32, Lemma 2.1]). Since a source of $H$ is a sink of $H^{-1}$, we deduce that $H$ has no sources as well. ∎
Let us resume the proof of Proposition 7.16. As $\tilde{h}$ is chain-transitive, by Lemma 7.18 we know that $\tilde{h}$ does not have sinks or sources. Therefore, as $\tilde{h}\in\mathcal{Z}$, Lemma 7.17 yields that $\tilde{h}$ is $\Omega-$stable. Thus, Theorem C ensures that $\tilde{h}$ is a topologically transitive Anosov diffeomorphism. Hence, by Corollary 7.2, for any pair of hyperbolic periodic points $p$ and $q$ of $\tilde{h}$ one has $\operatorname{ind}(p)=\operatorname{ind}(q)$. Yet, this contradicts the fact that $\tilde{h}$ has two hyperbolic periodic points $\tilde{p}_{1},\tilde{p}_{2}$ with different indices. Consequently, $f$ is $\Omega-$stable. Since $f$ is also chain-transitive, Theorem C implies that $f$ is a topologically transitive Anosov diffeomorphism. ∎
@@
When the dimension of the manifold is equal to $3$ or $4$ it is known that a $C^{1}-$robustly transitive diffeomorphism may not be an Anosov diffeomorphism [38, 54]. On the other hand, in [8], A. Arbieto and R. Ribeiro showed that the $C^{1}$ interior of the set of flows with the asymptotic average shadowing property (or with the average shadowing property) on a $3-$dimensional smooth closed manifold is formed by topologically transitive Anosov flows.
@@ 7.5. Application 5
$(b)\Rightarrow(c)$ This was established in [7, Proposition 3.1].
$(c)\Rightarrow(a)$ Take $f\in{\rm Diff}^{1}_{m}(M)$ complying with the condition $(c)$. We claim that $f$ is an Anosov diffeomorphism. Assume otherwise. By [7, Theorem 1.1], $f\notin\mathcal{F}^{1}_{m}(M)$, so there exists $g\in{\rm Diff}^{1}_{m}(M)$ close enough to $f$ with a non-hyperbolic periodic point $p$. Thus, according to [7, Lemma 5.1], there exists $h\in{\rm Diff}^{1}_{m}(M)$ which is close enough to $f$ and has two periodic points $q$ and $r$ with different indices. This contradicts $(c)$. ∎
@@
Denote by $\mathcal{KS}_{m}$ the $C^{1}$ generic set of Kupka-Smale volume-preserving diffeomorphisms provided by Robinson’s theorem (see [51]). The next lemma is a counterpart in the conservative setting of Corollary 7.3, and is proved by a similar argument.
@@
$(3)\Rightarrow(4)$ Suppose that, for any $\,p,q\in\operatorname{Per}_{h}(f)$, we have $\operatorname{ind}(p)=\operatorname{ind}(q)$. We claim that $f\in\mathcal{F}^{1}_{m}(M)$. Assume otherwise, that $f\notin\mathcal{F}^{1}_{m}(M)$, so there exists $g\in{\rm Diff}^{1}_{m}(M)$ close enough to $f$ with a non-hyperbolic periodic point $p$. Thus, according to [7, Lemma 5.1], there exists $h\in{\rm Diff}^{1}_{m}(M)$ which is close enough to $f$ and has two periodic points with different indices. Since $f\in\mathcal{J}$, by Lemma 7.22, $f$ has two hyperbolic periodic points with different indices. Yet, this contradicts the assumption on $f$. Therefore, $f\in\mathcal{F}^{1}_{m}(M)$.
$(4)\Rightarrow(5)$ This was established in [7, Theorem 1.1].
$(5)\Rightarrow(1)$ This is clear by Theorem A(c). ∎